# Mutual Information Free Topological Generalization Bounds via Stability

### v1

@@ 4 Empirical Evaluations
We conduct all experiments with a Vision Transformer (ViT) [DBK+21] on CIFAR-100 [KH+09]. Additional implementation details are left for the appendix.
@@
We follow the protocol of previous works [DDS23, ADS+24] and train models until convergence ($>93\%$ training accuracy), and use this checkpoint as a starting point for analysis of topological quantities. Our empirical analysis is then structured over a $4\times 4\times 5$ grid based on the following free variables: the number of training samples $n\in\{100,500,1000,5000,10000\}$, learning rates $\eta\in[10^{-6},10^{-4}]$, and batch sizes $b\in\{32,64,128,256\}$. We further investigated the stability parameter using a fixed learning rate of $10^{-6}$ and a batch size of $64$, testing across 10 different random seeds. These analyses are conducted without any further regularization techniques or learning rate scheduling to minimize the effect of confounding factors. All models are implemented in pytorch and trained with the ADAM optimizer.
Generalization error
As our theoretical results (see Section 3) provide bounds on the worst-case generalization gap $G_{S}(\mathcal{W}_{S,U})$, we seek to approximate this gap in practical settings. Prior work has typically used one of two proxies: the generalization gap at the final iteration, computed as final test risk -- final train risk [BLGS21], or the worst observed test risk over time, computed as worst-case test risk -- final train risk [ADS+24], where test risk was logged every $100$ iterations. To improve granularity, we assess both train and test risk every iteration and report worst-case (test risk - train risk) over all iterations.
Computing topological complexity
We employ the Krylov subspace approximation method to compute $\mathbf{PMag}(\mathcal{W}_{S,U})$, utilizing the preconditioned conjugate gradient (PCG) algorithm [Sal21]. Our computation of the $\mathbf{E}^{\alpha}(\mathcal{W}_{S,U})$ is based on the giotto-ph library [PHL+21], following the implementation of [ADS+24], using $\alpha=1$, which they argue is the most empirically relevant. In contrast to previous works [BLGS21, DDS23, ADS+24], which focus on detailed correlation analysis, our goal is to investigate how different topological quantities behave as a function of the sample size $n$, and how they relate instead of correlate to generalization gap. Hence, we refrain from introducing the granulated Kendall coefficients [JNM+19] and employ Pearson correlation coefficients to capture linear relationships. We present further correlation analysis in the appendix.
Distance matrix
The topological complexities that we consider in Section 3 are based on the Euclidean metric over the weight trajectory. Given a trajectory $\mathcal{W}_{S,U}$, the estimation of these quantities requires computing the distance matrix on all $\mathcal{W}_{S,U}$. To make this tractable, we uniformly sample $1500$ points from the full training trajectory, which consists of $5000$ iterations. Based on this sampled trajectory, we define the distance matrix using the Euclidean norm $\rho(w,w^{\prime}):=\left\|w-w^{\prime}\right\|$.
@@ 4.1 Results and Discussion
While previous works on topological generalization bounds [BLGS21, DDS23, ADS+24] analyzed the correlation between generalization and topological complexities, the intractability of IT components has thus far prevented a more complete analysis of the bounds. As our bound does not contain IT terms, relying primarily on the stability parameter defined by 3.1 and the topological quantities introduced in Theorems 3.5 and 3.6, all the terms they contain are amenable to numerical estimation, which is a unique feature among topological bounds. Next, we empirically characterize the interplay between stability and the topological complexities $\mathbf{E}^{\alpha}$ and $\mathbf{PMag}$.
Figure 1: Trajectory stability of the ViT for different training sample sizes $n$, for models with and without unseen injected samples $J$, for varying random seeds. The plot depicts models trained from a random initialization and converged pretrained models. We experiment with $\mathbf{Z}$ containing either validation or training points.
Stability Analysis
Our theoretical results depend on 3.1, which involves the stability of the trajectories $\mathcal{W}_{S,U}$ and provides a noisy variant of Definition 2.2. In our experiments, we assessed the assumption by a direct estimation of the stability coefficient. We focus on a condition similar to Definition 2.2 to avoid technical complications. Note that assuming the latter almost surely over $U$ is stronger than 3.1 so that the experiments can serve as an upper bound of Theorems 3.5 and 3.6.
To estimate the trajectory stability of each network, we replace a subset $\mathbf{J}$ of unseen samples for each training set $S\in\mathcal{Z}^{n}$ Section 4, where $|\mathbf{J}|=50$. After retraining the model with these perturbed datasets, we analyze the resulting trajectories $\mathcal{W}_{S,U}$ (original) and $\mathcal{W}_{S^{\prime},U}$ (perturbed). We quantify stability by computing the worst-case deviation in model losses across all iterations, analogous to a Haussdorf distance across the trajectories: $\max_{w\in\mathcal{W}_{S,U}}\min_{w^{\prime}\in\mathcal{W}_{S^{\prime},U}}\sup_{Z\in\mathbf{Z}}|\ell(w,Z)-\ell(w^{\prime},Z)|$, with $M:=\min\{n,500\}$ and $\mathbf{Z}:=\{Z_{1},\dots,Z_{M}\}$ a batch disjoint from $\mathbf{J}$, containing points either from the training or validation sets. Figure 1 shows the results for different sample sizes, where other hyperparameters are fixed. Note that we consider the loss difference over a batch $\mathbf{Z}$ of $m$ samples instead of taking the supremum over the whole dataset as in Definition 2.2. This procedures forms a tractable empirical substitute for Definition 2.2. To contextualize with 3.1 we report both the metric difference for train and validation sets.
The results are depicted in Figure 1. With increasing $n$, and thus, a decreasing proportion of the replaced samples, we observe that the per-iteration loss difference decreases, which is in accordance with established theoretical results [HRS16, BFGT20] and Lemma 3.2. Importantly, we observe a stable behavior when computing the loss differences for both the training and validation sets, suggesting that trajectory stability could be attained in practical settings with non-convex objectives.
This experiment shows that trajectory stability can be expected to hold in such settings. This observation might serve as a basis to understand the empirical success of previous studies [BLGS21, DDS23, ADS+24], in which the presence of the IT terms induced a discrepancy between theory and experiments. Our framework bridges this gap by rendering the bounds tractable. Moreover, Figure 1 suggests that stability holds when the experiment is started from a pre-trained weight, supporting previous findings that topological complexities correlate with $G_{S}(\mathcal{W}_{S,U})$ near convergence. ve that when the model is pretrained to full convergence, the per-iteration loss differences reflect a built-in stability, more so than for randomly initialized models.
Topological Complexity Analysis
Figure 2: Evolution of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ and $\mathbf{PMag}(100)$ with increasing sample size $n$, for learning rates $\eta=10^{-4}$ and $\eta=10^{-5}$.
@@
Figure 3: Variation of $\mathbf{E}^{1}$ with sample size $n$. Pearson correlation coefficients $r$ are reported for each subgroup.
Our stability analysis shows that our assumptions hold both for large sample sizes $n$ as well as for randomly initialized and converged models. While the stability parameter is expected to decrease with increasing sample size, the behavior of the topological complexity as a function of $n$ has not been empirically studied in the literature. We analyze the behavior of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ and $\mathbf{PMag}(100\cdot\mathcal{W}_{S,U})$ in Figure 3. While our theory suggests letting the scale in $\mathbf{PMag}$ vary with $n$, we fix the scale to $100$, which is on the order of magnitude suggested by the theory. This allows us to isolate the effect that $n$ has on $\mathcal{W}_{S,U}$ when observing the variations of the reported quantity. We report these quantities for fixed learning rates to avoid confounding with large differences in generalization. We observe that, with increasing sample size $n$, the values of $\mathbf{E}^{1}$ and $\mathbf{PMag}$ increase and seem to stabilize for large values of $n$. This interesting observation leads to an interdependence between the sample size and the topology of the training trajectory, a notion that has not been reflected in prior works. We leave a detailed theoretical investigation as future work.
The behavior of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ as a function of $n$ might be puzzling at first, as we expect the generalization error and, hence, our bound to decrease with $n$. There is no contradiction here as the topological complexities $\mathbf{C}(\mathcal{W}_{S,U})$ should increase at most polynomially with $n$. Indeed, the positive magnitude is always bounded in terms of the number of iterations, and, based on persistent homology dimension and limiting arguments as $n\to\infty$, $\mathbf{E}^{\alpha}(\mathcal{W}_{S,U})$ also has sub-polynomial growth as $n\to\infty$ under mild assumptions (e.g., bounded trajectories), see for instance [BLGS21]. As argued in Section 3, the stability coefficient $\beta_{n}$ decreases polynomially in many cases. As Theorems 3.5 and 3.6 are logarithmic in $\mathbf{C}(\mathcal{W}_{S,U})$, this ensures that the bounds cannot be vacuous as $n\to\infty$.
@@
In Figure 3, we observe that the sensitivity of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ with respect to the generalization gap (i.e., the slope of the affine regression of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ by the value of $G_{S}(\mathcal{W}_{S,U})$) increases when $n$ gets larger, which is what is predicted by our theory. Indeed, our main theorems assert that $\log\mathbf{E}^{1}(\mathcal{W}_{S,U})$ should be (approximately) of order at least $\beta_{n}^{-1/3}G_{S}(\mathcal{W}_{S,U})$, which amounts to $n^{1/3}G_{S}(\mathcal{W}_{S,U})$ in the event that $\beta_{n}=\Theta(1/n)$. Therefore, our experimental results strongly support Theorem 3.5.
Together, these observations unveil a strong coupling between stability and the topological complexity of the training trajectories. Moreover, the topological complexities increase for large $n$ and therefore become more relevant to the bound. The empirically observed coupling is supported by the theoretical results of Section 3, through the product of the stability parameter $\beta_{n}$ and the topological complexities.
@@ 5 Concluding Remarks
In this work, we present novel topological generalization bounds by making an assumption on the stability of training trajectories, demonstrating that this condition holds under practical settings. Based on this new framework, we obtain the first mutual information-free topological generalization bounds, which are, for the first time, tractable. This addresses one of the key theoretical issues in an emerging area of research. We empirically evaluate the interaction between the topological complexities, unveiling new connections between the sample size and the topology of the training trajectories. Our results show a strong coupling between stability and topological complexity and reveal compelling new behavior for varying sample size $n$. Empirical evidence supports our theoretical findings, suggesting a nuanced interplay between the stability-related and topological quantities that bound neural network generalization.
@@
While our bounds are more tractable than previous approaches with IT terms, $\beta_{n}$ is still challenging to quantify in the general case. Exploring suitable strategies to approximate this term poses exciting opportunities. On the theoretical side, the topological complexities based on the data-dependent pseudometric used in [ADS+24] are still not compatible with our approach and incorporating them is an object of future studies.
Broader impact
Due to its generality, our framework of provable generalization bounds based on topological quantities and trajectory stability is not anticipated to have any specific negative impact.
@@ Acknowledgements
The authors are grateful for the support of the Excellence Strategy of local and state governments, including computational resources of the LRZ AI service infrastructure provided by the Leibniz Supercomputing Center (LRZ), the German Federal Ministry of Education and Research (BMBF), and the Bavarian State Ministry of Science and the Arts (StMWK). T. B. was supported by a UKRI Future Leaders Fellowship (MR/Y018818/1) as well as a Royal Society Research Grant (RG/R1/241402). U.Ş. is partially supported by the French government under the management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute). B.D. and U.Ş. are partially supported by the European Research Council Starting Grant DYNASTY – 101039676.
Appendix A Trajectory Stability of Projected SGD on Nonconvex Problems
In Section 4.1, we empirically demonstrate that 3.1 is practical for modern neural networks like the Vision Transformer. What follows is a formal result demonstrating that 3.1 holds generally for projected Stochastic Gradient Descent (SGD) under a gradient Lipschitz condition.
Let us consider the projected SGD recursion on a loss function $\ell$, which can be non-convex. In particular, given a dataset $S=(z_{1},\dots,z_{n})\in\mathcal{Z}^{n}$, consider the following recursion:
$w_{k+1}=\Pi_{R}\left[w_{k}-\eta_{k}\nabla\ell(w_{k},z_{i_{k+1}})\right],$ (9)
where $i_{k}\in\{1,\dots,n\}$ is an index chosen at random at each iteration $k$ and $\Pi_{R}$ denotes the projection operator on the ball centered around 00 with radius $R$. Here, we fix the batch size to $1$ for clarity.
The next result establishes the trajectory-stability of SGD on Lipschitz and smooth (i.e., gradient Lipschitz) loss functions and is an adaptation of [HRS16, Theorem 3.12] to our setting.
Corollary A.1.
Fix $T\in\mathbb{N}$. Suppose that Assumption 3.3 holds, $\ell$ is gradient Lipschitz:
$\|\nabla\ell(w,z)-\nabla\ell(w^{\prime},z)\|\leq G\|w-w^{\prime}\|,$
for some $G<\infty$ and the step-size satisfies $\eta_{k}\leq c/k$ for some $c<1/G$. Then the SGD recursion on $K$ iterations in (9) is trajectory-stable with the stability parameter:
$\beta_{n}=\frac{4LR}{n-1}\left(\frac{L}{GR}\right)^{\frac{1}{Gc+1}}\sum_{k=1}^{T}k^{\frac{Gc}{Gc+1}}.$

### v2

@@ 5 Empirical Evaluations
In Section˜4, we provided the first fully computable topological/worst-case generalization bounds, without the need for intractable mutual information terms, which limited the empirical analysis of previous works.
Instead, our bounds are mainly based on the stability parameter $\beta_{n}$, which allows us to assess how they empirically compare with the worst-case generalization error and whether their structure aligns with observations, thus validating the framework and its practical applicability.
All experiments are performed using a Vision Transformer (ViT) (Dosovitskiy et al., 2021) on the CIFAR-100 dataset and GraphSAGE (Hamilton et al., 2017) on the MNISTSuperpixels dataset (Monti et al., 2017). All models are implemented in PyTorch and trained with the ADAM optimizer (Kingma and Ba, 2014). Further implementation details for both models are provided in Appendix˜C.
@@
We follow the protocol of previous works Dupuis et al. (2023); Andreeva et al. (2024) and train models until convergence, using this checkpoint as a starting point for further analysis. Our empirical analysis is structured in two main parts.
First, we analyze the order of magnitude of our bounds, offering initial insights into the tightness of our framework. We fine-tune each model for 500 iterations on a previously unseen dataset, with fixed batch sizes ($b$) and learning rates $(\eta)$, in particular $b\in\{64,128\}$ and $\eta\in\{10^{-6},10^{-5}\}$. For each $(b,\eta)$ configuration, we estimate the stability coefficient ˜3.1 over 5 different seeds.
Then we examine the relationship between stability and topological complexity (i.e., weighted life-time sums and positive magnitude in Theorem˜4.4). This analysis is motivated by the dependence of the stability parameter on the sample size $n$ and the multiplicative interaction between stability and topological complexity in Theorem˜4.4. To this end, we fine-tune the models for further $5\cdot 10^{3}$ iterations. The experiments are structured on a $4\times 4\times 5$ grid based on the learning rate $\eta$, the batch size $b$, and $n$. For detailed information about the hyperparameter, we refer to Appendix˜C.
Quantity Estimation
In this paragraph, we describe the quantities used in our empirical analysis and their estimation. Implementation details are provided in the Appendix.
  • Generalization error: We approximate $G_{S}(\mathcal{W}_{S,U})$ (Equation˜4) by tracking train and test risk at every iteration and reporting $\max_{t}\{\texttt{test risk}-\texttt{train risk}\}$. This refines prior proxies based on either the final gap (Birdal et al., 2021) or the worst test risk (Andreeva et al., 2024).
  • Topological complexities: We compute $\mathbf{PMag}(\mathcal{W}_{S,U})$ via a Krylov subspace PCG approximation Salim (2021), and $\mathbf{E}^{\alpha}(\mathcal{W}_{S,U})$ using giotto-ph Pérez et al. (2021) with $\alpha=1$.
  • Distance matrix: For trajectories $\mathcal{W}_{S,U}$, we form a Euclidean distance matrix $\rho(w,w^{\prime})=\|w-w^{\prime}\|$. To reduce computational cost, we uniformly sample $1500$ of the $5000$ iterations.
  • Stability parameter: For random set stability (˜3.1), we replace $50$ unseen samples per training set, retrain to obtain $\mathcal{W}_{S,U}$ and $\mathcal{W}_{S^{\prime},U}$, and measure worst-case loss deviations across iterations $\max_{w\in\mathcal{W}_{S,U}}\;\min_{w^{\prime}\in\mathcal{W}_{S^{\prime},U}}\;\sup_{Z\in\mathcal{Z}}\;\big|\ell(w,Z)-\ell(w^{\prime},Z)\big|$. With $M=500$ held-out points $\mathbf{Z}$ (see Algorithm˜1), we average results over $5$ different random seeds. Note that this method necessarily leads to an optimistic estimation of the stability parameter $\beta_{n}$, as it would be intractable to evaluate the supremum over the entire data space $\mathcal{Z}$.
@@ 5.1 Results and Discussion
We now present our main empirical results. First, we numerically evaluate our bounds. Next, we investigate the interplay between stability and the topological quantities appearing in Theorem˜4.4.
Order of the bounds
Our goal is to have an estimate of the order of magnitude of the worst-case error predicted by our bounds. To avoid the computationally costly evaluation of Lipschitz constants, we estimate a simple upper bound on the Rademacher complexity that is common to all our theoretical results. Concretely, we use Massart’s lemma (Shalev-Shwartz and Ben-David, 2014) to bound the right-hand side of Equation˜8 by $2\sqrt{2\log(T)/J}+2J\beta_{n}$, where $T$ is the number of iterations. We estimate $\beta_{n}$ using the procedure above and optimize over $J$ (see Section˜C.3).
We present the resulting estimation of our bound in Table˜1, for different learning rates and batch sizes. A consistent pattern can be observed: the estimated bounds are typically close to an order of magnitude larger than the actual worst-case generalization error. However, in most experimental settings, the estimated bounds remain below $100\%$ accuracy, hence, provide meaningful guarantees. We also find that the stability parameter $\beta_{n}$ varies with the hyperparameters $(\eta,b)$ and, therefore, encodes the impact of the learning algorithm on generalization error.
Moreover, the estimated $\beta_{n}$ captures the generalization behavior: specifically, $G_{S}(\mathcal{W}_{S,U})$ tends to decrease as $\beta_{n}$ becomes smaller, a trend that holds for both ViT and GraphSage.
Table 1: Comparison of ViT and GraphSage across different learning rates ($\eta$) and batch sizes ($b$). We report $\beta_{n}$, $G_{S}(W_{S,U})$, and the resulting bound. We use the 0-1 loss.
Model $\bm{\eta}$ $\bm{b}$ $\bm{10^{4}\cdot\beta_{n}}$ $\bm{10^{2}\cdot G_{S}(W_{S,U})}$ $\bm{10^{2}\cdot}$ Bound ViT $10^{-4}$ 64 4.72 $\pm$ 0.52 10.24 $\pm$ 0.36 104.43 $10^{-4}$ 128 5.52 $\pm$ 0.67 12.84 $\pm$ 0.55 105.24 $10^{-5}$ 64 2.16 $\pm$ 0.36 7.16 $\pm$ 0.33 68.47 $10^{-5}$ 128 2.24 $\pm$ 0.22 6.36 $\pm$ 0.33 68.67 GraphSage $10^{-4}$ 64 2.56 $\pm$ 0.36 5.48 $\pm$ 0.36 75.63 $10^{-4}$ 128 2.64 $\pm$ 0.22 5.20 $\pm$ 0.40 75.79 $10^{-5}$ 64 0.64 $\pm$ 0.22 4.60 $\pm$ 0.14 47.79 $10^{-5}$ 128 0.80 $\pm$ 0.00 4.84 $\pm$ 0.17 48.59
To the best of our knowledge, we are the first to fully estimate a bound on the worst-case error, thereby providing a stronger guarantee than in prior work. In contrast, previous studies focused on bounds for a single weight vector and similarly reported discrepancies of one to two orders of magnitude between the estimated bound and the actual error (Chen et al., 2018; Li et al., 2019; Dupuis et al., 2023; Harel et al., 2025). This suggests that our worst-case generalization bounds are reasonable tight. Moreover, the fact that smaller estimated generalization gaps are consistently associated with smaller stability parameters and, therefore, smaller bounds, shows that the bounds adapt meaningfully to model performance.
@@
Figure 2: Variation of $\mathbf{E}^{1}$ with sample size $n$ for the model ViT. Pearson correlation coefficients $r$ are reported for each subgroup.
Figure 3: Variation of $\mathbf{E}^{1}$ with sample size $n$ for the model GraphSage. Pearson correlation coefficients $r$ are reported for each subgroup.
Interplay between Stability and $\mathbf{C}(\mathcal{W}_{S,U})$
@@
In Figure˜3 and 3, we observe that the sensitivity of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ with respect to the generalization gap (i.e., the slope of the affine regression of $\mathbf{E}^{1}(\mathcal{W}_{S,U})$ by the value of $G_{S}(\mathcal{W}_{S,U})$) increases when $n$ gets larger, which is what is predicted by our theory. Indeed, Theorem˜4.4 assert that $\log\mathbf{E}^{1}(\mathcal{W}_{S,U})$ should be (approximately) of order at least $\beta_{n}^{-1/3}G_{S}(\mathcal{W}_{S,U})$, which amounts to $n^{1/3}G_{S}(\mathcal{W}_{S,U})$ in the event that $\beta_{n}=\Theta(1/n)$. Therefore, our experimental results strongly support Theorem˜4.4. Together, these observations unveil a strong coupling between stability and the topological complexity of the training trajectories. Moreover, the topological complexities increase for large $n$ and therefore become more relevant to the bound. The empirically observed coupling is supported by the theoretical results of Section˜3, through the product of the stability parameter $\beta_{n}$ and the topological complexities. Finally, we observe that the reported Pearson correlations are lower for bigger $n$, especially for the GraphSage experiment. We believe that it could be due to the fact that reaching local minima is harder when $n$ increases, which can decrease the correlation between the topological complexities and the generalization error, as it was already reported in (Birdal et al., 2021; Andreeva et al., 2024).
@@ 6 Concluding Remarks
In this work, we present a new framework for obtaining worst-case generalization bounds over data-dependent random sets. This framework is based on a new concept of random set stability, which we demonstrate holds for practically used algorithms, and is related to classical stability notions. We then bound the worst-case generalization error by the (weighted) sum of the random set stability parameter and a Rademacher complexity term. We present several applications of our new techniques by recovering classical generalization bounds and proving mutual information-free versions of existing fractal and topological bounds, which makes them for the first time fully computable. We supported our theory through a series of experiments conducted in practically relevant settings.
@@
Despite providing the first mutual information-free topological and fractal bounds, our main results have two main disadvantages: (i) we only provide expected bounds, and (ii) our framework only allows for Euclidean-based topological complexities, and we typically do not cover the case of the data-dependent pseudometric introduced by Dupuis et al. (2023). Other methods to remove information-theoretic terms exist in the literature. For instance, Neu et al. (2021) added Gaussian noise to the algorithm output in their study of the generalization error of SGD. Extending such a technique to the random set setting is a promising direction for future works. Finally, while our paper extends uniform stability to random sets, we believe that other notions of stability, like the one of Lei and Ying (2020) could be extended to our random set setting and might lead to better generalization bounds under specific assumptions.
Ethics statement
This work adheres to the Code of Ethics, ensuring responsible use of publicly available datasets and maintaining full transparency in experimental design and reporting. The study complies with ethical guidelines on the integrity of research and considers possible social impacts. We have not used data collected on human subjects. In order to minimize carbon emissions, we utilized high-end GPUs only when they are necessary to train our large models. Since contributions are primarily theoretical, the focus is on advancing our understanding of the theory of artificial intelligence rather than targeting specific applications that may have direct societal impacts.
Reproducibility Statement
All data sets used in this work are referenced and publicly accessible. The experimental configurations and computational environment are outlined in detail within the main text and the appendix. The original code, accompanied by instructions, is included in the Supplementary Material. Lastly, all random seeds are provided to facilitate precise reproducibility. We plan to make our implementation available upon publication.
@@ Acknowledgements
The authors are grateful for the support of the Excellence Strategy of local and state governments, including computational resources of the LRZ AI service infrastructure provided by the Leibniz Supercomputing Center (LRZ), the German Federal Ministry of Education and Research (BMBF), and the Bavarian State Ministry of Science and the Arts (StMWK). T. B. was supported by a UKRI Future Leaders Fellowship (MR/Y018818/1) as well as a Royal Society Research Grant (RG/R1/241402). U.Ş. is partially supported by the French government under the management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute). M.T, B.D. and U.Ş. are partially supported by the European Research Council Starting Grant DYNASTY – 101039676.
Appendices
We now provide additional technical details and proofs that are omitted from the main write-up. Further, we provide additional technical and empirical results. The appendix is thus organized in the following way:
  • Appendix A provides supplementary technical background.
  • Appendix B contains the omitted proofs of all our theoretical results.
  • Appendix˜C details the experimental setup and procedures required for reproducibility.
  • Appendix˜D presents additional empirical results.
Appendix A Additional technical background
In this section, we provide the relevant definitions as well as the technical background for our results.
A.1 Rademacher Complexity
Recent advances on topological generalization bounds are based on a data-dependent Rademacher complexity, leveraging PAC-Bayesian theory on random sets Dupuis et al. (2024). In our stability-based approach, the Rademacher complexity also plays a central role. We define the standard Rademacher complexity Shalev-Shwartz and Ben-David (2014); Bartlett and Mendelson (2002) as:
Definition A.1.
Let $\mathcal{W}\subset\mathbb{R}^{d}$, $S\in\mathcal{Z}^{n}$, and $\bm{\epsilon}:=(\epsilon_{1},\dots,\epsilon_{n})\sim\mathcal{U}(\{-1,1\})^{\otimes n}$ be i.i.d Rademacher random variables (i.e., centered Bernoulli variables). Let $\ell:\mathcal{W}\times\mathcal{Z}\to\mathbb{R}_{+}$ be a loss function. We define the Rademacher complexity of $\mathcal{W}$ (relative to $\ell$) as:
$\mathbf{Rad}_{S}(\mathcal{W}):=\mathbb{E}_{\bm{\epsilon}}\left[\sup_{w\in\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\ell(w,z_{i})\right].$
A.2 Covering and Packing Numbers
In this section, we provide definitions of covering and packing numbers. These quantities have long been central in learning theory, particularly in the context of classical covering-based arguments for Rademacher complexity Shalev-Shwartz and Ben-David (2014); Bartlett and Mendelson (2002).
Definition A.2 (Covering number).
Let $(\mathcal{W},\rho)$ be a metric space with $\mathcal{W}\subset\mathbb{R}^{d}$ of finite diameter, and let $\delta>0$. The covering number $N_{\delta}(\mathcal{W})$ is the smallest integer $K\in\mathbb{N}^{\star}$ such that there exist points $x_{1},\dots,x_{K}\in\mathcal{W}$ satisfying
$\mathcal{W}\subset\bigcup_{k=1}^{K}B_{\rho}(x_{k},\delta),$
where $B_{\rho}(x,\delta)$ denotes the open ball of radius $\delta$ centered at $x$ with respect to the metric $\rho$.
In this definition, we take the centers $(x_{1},\dots,x_{K})$ of the covering to be in $\mathcal{W}$. This allows us to leverage the "local" Lipschitz continuity assumption defined in ˜4.1. This definition does not lose any generality compared to taking the points in $\mathbb{R}^{d}$. In particular, both definitions yield the same fractal dimensions (Falconer, 2013).
Definition A.3 (Packing number).
Let $(\mathcal{W},\rho)$ be a metric space and $\delta>0$. The packing number $P_{\delta}^{\rho}(\mathcal{W})$ is the cardinality of a maximal set of points in $\mathcal{W}$ such that the closed balls of radius $\delta$ centered at these points are pairwise disjoint.
A.3 Intrinsic Dimensions
In this section, we focus on the upper box-counting dimension (also known as the upper Minkowski dimension), one of the most widely used notions of fractal dimension Falconer (2013); Mattila (1999). This notion has also found applications in machine learning, for instance in the works of Simsekli et al. (2020); Dupuis et al. (2023).
Definition A.4 (Upper box-counting dimension).
Let $(\mathcal{W},\rho)$ be a metric space of finite diameter and let $(|N_{\delta}(\mathcal{W})|)_{\delta>0}$ denote the covering numbers introduced in Definition˜A.2. The upper box-counting dimension is defined by:
$\overline{\dim}_{\mathrm{B}}(\mathcal{W}):=\limsup_{\delta\to 0}\frac{\log|N_{\delta}(\mathcal{W})|}{\log(1/\delta)}.$
We refer the reader to (Falconer, 2013) for more technical background on this notion.
Remark A.1.
It has been shown in Kozma et al. (2006); Schweinhart (2020) that, for any compact metric space, several notions of dimension coincide with the well-known upper box-counting dimension (see Definition˜A.4). In particular, the celebrated persistent homology dimension Adams et al. (2020) has been connected to generalization Birdal et al. (2021) and has also been studied empirically in various works Andreeva et al. (2024); Tan et al. (2024).
A.4 Topological Complexities
In this section, we define the topological quantities appearing in Theorem˜4.4.
$\alpha$-weighted lifetime sum ($\mathbf{E}^{\alpha}$)
The weighted lifetime sums is a central notion in the study of persistent homology, a subfield of topological data analysis (TDA). The interested reader may find a full exposition of these notions in (Boissonnat et al., 2018; Schweinhart, 2020; Kozma et al., 2006). There exist several equivalent definitions the weighted lifetime sums of degree 0, see (Andreeva et al., 2024) for a concise review of these definitions in the context of machine learning.
In this section, we present an intuitive approach to the weighted lifetime sum, which we also restrict to finite subsets of $\mathbb{R}^{d}$ and to homology of dimension 0. Note that this is a simplified, non-standard (though equivalent) definition of $\mathrm{PH}^{0}$.
Definition A.5 (Persistent homology of degree 0 $\mathrm{PH}^{0}$).
Let $\mathcal{W}\subset\mathbb{R}^{d}$ be a finite set with cardinality $N$, endowed with the Euclidean metric. For each $t\geq 0$ (which is often referred to as a “time” variable in the TDA literature), we construct an undirected graph $G_{t}$ with edges given by
$\forall x,y\in\mathcal{W},\quad\{x,y\}\in G_{t}\iff\rho(x,y)\leq t.$
There exists a finite sequence of times $0<t_{1}<\dots<t_{k}<+\infty$ such that the number of connected components in $G_{t_{i}}$ changes compared to $G_{t}$ for $t<t_{i}$. Let $c_{i}$ denote the number of connected components in $G_{t_{i}}$. By convention, set $c_{0}=N$ and $t_{0}=0$, and define $n_{i}:=c_{i}-c_{i-1}.$ Then $\mathrm{PH}^{0}$ is defined as the multi-set,
$\mathrm{PH}^{0}(\mathcal{W}):=\Bigl\{\!\!\bigl\{\underbrace{t_{1},\dots,t_{1}}_{n_{1}\text{ times}},t_{2}\dots,\underbrace{t_{k},\dots,t_{k}}_{n_{k}\text{ times}}\bigr\}\!\!\Bigr\},$
where the notation $\{\!\{\cdot\}\!\}$ denotes a multi-set.
Remark A.2.
The construction of the graph above consists in looking at the $1$-dimensional simplices of the Rips complex associated with $\mathcal{W}$, see (Boissonnat et al., 2018, Chapter 2).
We now use $\mathrm{PH}^{0}$ to define the key quantities in our work, in particular the $\alpha$-weighted sum $\mathbf{E}_{\rho}^{\alpha}$.
Definition A.6 ($\alpha$-weighted lifetime sums).
Let $\mathcal{W}\subset\mathbb{R}^{d}$ be a finite subsets of $\mathbb{R}^{d}$, endowed with the Euclidean metric. For $\alpha\geq 0$, the $\alpha$-weighted lifetime sum of $\mathcal{W}$ is defined as
$\mathbf{E}^{\alpha}(\mathcal{W}):=\sum_{t\in\mathrm{PH}^{0}(\mathcal{W})}t^{\alpha},$
where $\mathrm{PH}^{0}$ denotes the set of lifetimes of the 0-dimensional persistent homology intervals of $\mathcal{W}$, or, equivalently, the above definition.
Magnitude and positive magnitude
Magnitude is another known topological invariant used in TDA, it was introduced by Leinster (2013). In the context of machine learning, it has been used in particular by Andreeva et al. (2023) to analyze the generalization error of neural networks.
A detailed account of this theory can be found in (Meckes, 2013, 2015). In this paragraph, we restrict ourselves to a particular case, that is enough for our purpose. In particular, we only define these notions for finite sets, but all our theoretical results can be seamlessly extended to compact subsets of $\mathbb{R}^{d}$ (see (Andreeva et al., 2024, Remark B.15)).
In all this paragraph, we fix a finite set $\mathcal{W}\subset\mathbb{R}^{d}$. The key notion behind both magnitude and positive magnitude is that of a weighting of $\mathcal{W}$, which is defined below.
Definition A.7 (Weighting).
Let $\lambda>0$. A $\lambda$-weighting of $\mathcal{W}$ is a vector $\beta_{\lambda}\in\mathbb{R}^{\mathcal{W}}$ such that:
$\forall a\in\mathcal{W},~\sum_{b\in\mathcal{W}}e^{-\lambda\left\|a-b\right\|}\beta_{\lambda}(b)=1.$
The existence of such a weighting has been shown by Leinster (2013).
We can now define the magnitude of $\mathcal{W}$.
Definition A.8 (Magnitude, (Leinster, 2013)).
With the notations of Definition˜A.7, the magnitude function of $\mathcal{W}$ is defined as:
$(0,+\infty)\ni\lambda\longmapsto\mathbf{Mag}(\lambda\cdot\mathcal{W}):=\sum_{a\in\mathcal{W}}\beta_{\lambda}(a).$
Recently, Andreeva et al. (2024) demonstrated that by slightly adapting the definition of magnitude, one obtains an object that is naturally related to the Rademacher complexity (Definition˜A.1), hence, relating it to learning theory. Therefore, they introduced the notion of positive magnitude, which we recall below.
Definition A.9 (Positive magnitude, (Andreeva et al., 2024)).
With the notations of Definition˜A.7, the positive magnitude function of $\mathcal{W}$ is defined as:
$(0,+\infty)\ni\lambda\longmapsto\mathbf{PMag}(\lambda\cdot\mathcal{W}):=\sum_{a\in\mathcal{W}}\left(\beta_{\lambda}\right)_{+}(a),$
where $x_{+}:=\max(x,0)$.
Appendix B Omitted Proofs
In this section, we present the omitted proofs in full detail. We also present a few additional technical results that are useful for our proofs and may be of independent interest.
B.1 Proof of Corollary˜3.3
We give below the proof of Corollary˜3.3, which provides a random set stability assumption for projected SGD, in smooth and Lipschitz continuous settings.
See 3.3